# Timing performance of SiPM-on-Tile elements: Laboratory and test beam measurements

### v1

@@ Abstract
The SiPM-on-Tile technology for highly granular calorimeters, where small plastic scintillator tiles are directly read out with SiPMs, has been developed for the CALICE Analog Hadron Calorimeter, and has been adopted for parts of the hadronic section of the CMS High Granularity Calorimeter. For future electron-positron colliders, a single cell time stamping on the sub-nanosecond level for energy deposits corresponding to single minimum-ionizing particles is desired to provide background rejection and to support pattern recognition and energy reconstruction with particle flow algorithms. To better understand the intrinsic time resolution achievable with the SiPM-on-Tile technology, we have performed detailed measurements in test beams at DESY, probing scintillator tiles of different sizes. The study is complemented by laser measurements that provide insights into processes within the scintillator tile relevant for timing. In this publication, we will discuss our measurement methods and the results of our SiPM-on-Tile timing study.
@@ 1 Introduction
Timing in event and shower reconstruction has become one of the main directions in calorimeter development and shows promising potential for the improvement of energy resolution in highly granular hadronic calorimeters [1, 2, 3]. High-precision time information allows to characterise the evolution of particle showers with the aim to identify shower components. Other uses of timing in calorimetry include background and pileup suppression, as well as the search for new long-lived particles.
@@
The SiPM-on-Tile technology has been developed for the CALICE Analog Hadron Calorimeter (AHCAL) [5], and has been adopted for parts of the hadronic section of the CMS High Granularity Calorimeter (HGCAL) [6]. The plastics scintillator tiles are square-shaped, wrapped in reflective foil, and are directly placed on a PCB which hosts the SiPMs and the front-end electronics. A central dimple in each tile houses the SiPM which reads out the light from that tile directly. This configuration is optimised for light collection and allows automatic assembly with a glue dispenser and a pick-and-place machine.
To better understand the intrinsic time resolution achievable with the SiPM-on-tile technology, we have performed detailed measurements in beam tests at DESY, probing different scintillator tile sizes. In addition, we introduce measurement methods that allow to observe the time structure of the scintillation process, the time structure of the light collection in the scintillator tile, and the response of the SiPM and measurement electronics separately.
@@ 1.1 Experimental Equipment
Modules based on a single SiPM-on-tile element were developed for the CALICE T3B experiment [7]. These modules were further developed and used within the CLAWS project for beam background studies and a beam abort system for the SuperKEKB collider [8, 9, 10]. For the measurements discussed in this publication, we used the latest iteration of the CLAWS boards, the phase 3 CLAWS shown in figure 1. Each boards contains a single SiPM under a plastic scintillator tile, as well as the corresponding amplifier electronics allow an analog transmission of the sensor signals over extended distances.
@@
The analog signals are transmitted over CAT6a ethernet cables, which are also used to provide the supply voltage for the amplifier circuit and the bias voltage of the SiPM. The full signal waveforms of each event are recorded with a PicoScope 6804E digital computer-controlled oscilloscope [11].
The SiPMs installed on the CLAWS boards are Hamamatsu S13360-1325PE [12], the same that are used in the AHCAL technological prototype [5].
@@ 1.2 Breaking Down the Signal Creation
Figure 2 illustrates the three signal creation steps as well as the different measurement methods. The following section 2 shows our test beam measurements. Section 3 discusses the timing properties of BC408 measured with small scintillator cubes, and the laser measurements are presented in section 4.
@@ 2 Test Beam Measurements
Measurements with high-energy charged particles involve all three steps of signal creation (scintillation, light collection and light detection at the SiPM). We used electrons with an energy of 3 GeV at DESY’s test beam facility [15]. In two test beam weeks of the CALICE AHCAL group, in October 2020 and 2021, we placed our timing setup in front of the AHCAL prototype, such that both experiments used the same electron beam.
@@ 2.1 Experimental Setup
The measurements are controlled by a python script that regularly (approximately every 5 minutes) interrupts for calibration. A calibration run is a separate measurement where each channel is recorded individually, one after another. The trigger level is set to approximately 0.5 p.e which means that it triggers mostly on dark counts where a single pixel of the SiPM fired. Accordingly, the voltage range is set to ±$\pm 50$ mV. These calibration measurements allow the monitoring of the 1 p.e. signal, and with that the stability of the SiPM gain for each channel. The integral of the calibration waveforms is used to estimate the number of photoelectrons for the recorded signal waveforms.
@@
$20\times 20\times 3~{}\text{mm}^{3}$ $20\times 20\times 3~{}\text{mm}^{3}$$\pm 1$ V, 8 bit External October 2021
$30\times 30\times 3~{}\text{mm}^{3}$ $30\times 30\times 3~{}\text{mm}^{3}$$\pm 1$ V, 8 bit Internal October 2020
@@
Calibration Trigger on 1 p.e. peak ±$\pm 50$ mV, 8 bit Internal
@@ 2.2 Results: Light Yield
The light yield of a detector module is defined as the most probable number of photoelectrons measured at the SiPM for minimum ionization (1 MIP of energy deposition) in the scintillator tile. The light yield is the most probable signal area of a minimum ionizing particle signal divided by the area of a 1 p.e. signal.
@@
To extract the most probable energy deposition, we performed a fit to a Landau distribution convoluted with a Gaussian distribution, commonly referred to as "Langauss". The Gaussian part accounts for detector noise and measurement uncertainties. The energy deposition in the smallest $20\times 20\times 3~{}\text{mm}^{3}$ scintillator tiles (figure 4(a)) agrees very well with the expected Langauss behaviour.
The distributions for the larger scintillator tiles (figures 4(b) and 4(c)) on the other hand show an excess of events at energies above the minimum ionization peak [^2] Especially the run with the $40\times 40\times 3~{}\text{mm}^{3}$ scintillator tiles (figure 4(c)) has a high umber of multi-particle events, for once because larger tiles have a higher geometric cross section, and on the other hand because a beam collimator with a larger aperture was used during this measurement.. This can be attributed to events where more than one particle deposits energy in the scintillator tiles. To make sure that no double particle events are included in the analysis, an upper limit for the fit range was set manually before the fit, indicated by the black lines in figure 4. Effectively, these limits correspond to an energy of approximately 1.5 MIP.
@@ 2.3 Results: Time Resolution
Hit times outside of the interval ±$\pm 5~{}$ns are rejected since they tend to be due to early hits caused by baseline noise on one of the two channels.
@@
Table 3: Single channel time resolution at minimum ionization $\sigma_{1MIP}$ for different BC408 scintillator tiles measured at the test beam. The total uncertainty includes the uncertainty from the fit as well as expected fluctuations in the time resolution that arise from tile-to-tile differences.
Tile size $\sigma_{1MIP}$ (ns) Fit reduced $\chi^{2}$ Fit uncertainty (ns) Total uncertainty (ns)
$20\times 20\times 3~{}\text{mm}^{3}$ 0.3847 3.10 0.0004 0.023
$30\times 30\times 3~{}\text{mm}^{3}$ 0.5785 1.54 0.0006 0.035
$40\times 40\times 3~{}\text{mm}^{3}$ 0.7027 1.76 0.0008 0.042
@@ 3 Timing Properties of BC408
The timing properties of the scintillation light emission of the BC408 plastic scintillator are expected to have an important impact on the time resolution of the SiPM-on-tile module. The influence of the scintillator properties can be accessed experimentally by reducing geometric effects in the scintillator as far as possible.
3.1 Experimental Setup
For measuring the timing properties of BC408, we use a setup that is designed to avoid geometric effects as far as possible and only detects light directly emitted from the scintillator. Scintillator cubes measuring $5\times 5\times 5~{}\text{mm}^{3}$ are used instead of larger tiles to minimize the path that photons travel. Instead of ESR foil, the cubes are wrapped in black foil to suppress multiple reflections The scintillator cubes are excited using electrons from a 90Sr source. The signal of one cube is used for the timing analysis, and the second cube is used to provide a coincidence of two cubes as a trigger signal. This ensures full penetration of the first cube, and thus scintillation light emission over the full depth of the scintillator element. The measurement setup is illustrated in figure 7, and the measurement settings are summarised in table 4.
@@ 3.1 Experimental Setup
Since each photon triggers the same pixel response in the SiPM and the shape of this 1 p.e. signal is known from calibration measurements, the arrival times of single photons at the SiPM can be estimated from the analog waveform measured with the oscilloscope. To this end we have developed a waveform decomposition algorithm, inspired by the signal reconstruction in the T3B experiment [7]. Starting from the measured waveform, the algorithm iteratively subtracts 1 p.e. signals until only the baseline noise remains. The points in time where the 1 p.e. responses were subtracted are used as photon hit times. A detailed description of the waveform decomposition algorithm, as well as a discussion of its performance, can be found in [16].
@@
BC408 $5\times 5\times 5~{}\text{mm}^{3}$ cubes ±$\pm 500$ mV, 8 bit 0.4 ns Light from first BC408 cube
Laser measurements ±$\pm 1$ V, 8 bit 0.4 ns Laser driver
Figure 7: Setup with small scintillator cubes used to mitigate geometric effects when measuring the photon emission time spectrum. The scintillator cubes are made of BC408 and have a dimension of $5\times 5\times 5~{}\text{mm}^{3}$. A black foil shields the cubes from ambient stray light and crosstalk. The oscilloscope is configured as a coincidence trigger on both channels.
3.2 Results
The photon emission time distribution of BC408, measured with small scintillator cubes, is shown in figure 8. The red line is a fit to a bi-exponential function as in equation 3.1, convoluted with a Gaussian to account for detector and electronic noise.
$p(t|\tau_{rise},\tau_{fall})=\begin{cases}\exp\left(-\frac{t}{\tau_{fall}}\right)\cdot\left[1-\exp\left(-\frac{t}{\tau_{rise}}\right)\right]\cdot\frac{\tau_{rise}+\tau_{fall}}{\tau_{fall}^{2}}&t>0\\ 0&\text{else}\end{cases}$ (3.1)
Table 5 compares the rise and fall times of the BC408 photon emission spectrum, obtained with the fit in figure 8, to the rise and fall times from the BC408 datasheet [13]. Since the manufacturer does not state how their time constants are defined, these values are however not directly comparable.
Figure 8: Emission time spectrum of BC408 measured with small scintillator cubes. The photon hit times were extracted by decomposition of the waveform into 1 p.e. peaks (see section 3.1). The solid line is a fit to the bi-exponential function in equation 3.1 convoluted with a Gaussian.
Table 5: Timing properties of the plastic scintillator BC408 as measured in our setup, compared to the manufacturer information from the BC408 datasheet. Note that the numbers are not directly comparable.
Rise time (ns) Fall time (ns)
Measured 0.42 $\pm$ 0.18 2.61 $\pm$ 0.15
Manufacturer information 0.9 2.1
@@ 4.1 Experimental Setup
Figure 9 shows an illustration and a photograph of the laser setup. The light intensity that reaches the SiPMs is controlled using a set of neutral density optical filters and is monitored with a power meter. The beam is split and directed to two different SiPMs, which allows to calculate the time resolution from the hit time difference, as described in section 2.3. The trigger signal is directly provided from the laser driver.
@@
To probe the response of the SiPM and electronics to very short light pulses, the laser beam is aimed directly onto the SiPM. This allows to estimate the timing uncertainty that arises from the light sensor, the electronics and the data acquisition, i.e. the timing uncertainty of the light detection.
@@
4.2 Results: Light Detection and Hardware Response
Figure 10 shows the time resolution of the light detection which was obtained from laser measurements, compared to the time resolution of a $30\times 30\times 3~{}\text{mm}^{3}$ scintillator tile at the test beam. If we assume that the timing uncertainty from the electronics and SiPM is uncorrelated to the effects of the scintillator tile (scintillation and light collection), we can write the overall time resolution of the SiPM-on-tile system as
@@ 4.2 Results: Light Detection and Hardware Response
The green data points in figure 10 are the time resolution of the test beam measurement without the uncertainties induced by the time resolution of the SiPM:
@@
Figure 10: The energy dependent time resolution of a test beam measurement with BC408 $30\times 30\times 3~{}\text{mm}^{3}$ scintillator tiles is compared to the time resolution obtained from short laser pulses aimed directly at the SiPM, which is essentially the time resolution of the SiPM and electronics. The green curve is the computed time resolution of the test beam measurement without the timing uncertainties of the SiPM. The common energy scale is photoelectrons, and for reference the MIP scale of the test beam measurement is shown at the top axis.
@@
This can also be seen the time distribution of photons arriving at the SiPM, shown in figure 11. Compared to the time distribution of scintillation photons, the laser measurement yields a very short time distribution. Further discussion on this plot can be found in [16].
Figure 11: Time distribution of photoelectrons measured at the SiPM. The orange histogram is the time distribution of the SiPM and electronics, triggered by short laser pulses. The time distribution of scintillation photons from a test beam measurement is shown in blue. The photon hit times were extracted by decomposition of the waveform into 1 p.e. peaks (see section 3.1).
@@
All in all, the laser measurements show that the SiPM, CLAWS, amplifier electronics and oscilloscope together are still significantly faster than other signal parts. Therefore, the contribution of the hardware to the time structure of the measured signals can be neglected.
4.3 Results: Light Collection

### v2

@@ Abstract
The SiPM-on-Tile technology for highly granular calorimeters, where small plastic scintillator tiles are directly read out with SiPMs, has been developed for the CALICE Analog Hadron Calorimeter, and has been adopted for parts of the hadronic section of the CMS High Granularity Calorimeter. For future electron-positron colliders, a single cell time stamping on the sub-nanosecond level for energy deposits corresponding to single minimum-ionizing particles is desired to provide background rejection and to support pattern recognition and energy reconstruction with particle flow algorithms. To better understand the time resolution achievable with the SiPM-on-Tile technology, we have performed detailed measurements in test beams at DESY, probing scintillator tiles of different sizes. The study is complemented by laser measurements that provide insights into processes within the scintillator tile relevant for timing. In this publication, we will discuss our measurement methods and the results of our SiPM-on-Tile timing study.
@@ 1 Introduction
Timing in event and shower reconstruction has become one of the main directions in calorimeter development and shows promising potential for the improvement of energy resolution in highly granular hadronic calorimeters [1, 2, 3, 4]. High-precision time information enables the characterisation of the evolution of particle showers with the aim to identify shower components. Other uses of timing in calorimetry include background and pileup suppression, as well as the search for new long-lived particles.
@@
The SiPM-on-Tile technology has been developed for the CALICE Analog Hadron Calorimeter (AHCAL) [6], and has been adopted for parts of the hadronic section of the CMS High Granularity Calorimeter (HGCAL) [7]. The plastic scintillator tiles are square-shaped, wrapped in reflective foil, and are directly placed on a PCB which hosts the SiPMs and the front-end electronics. A central dimple in each tile houses the SiPM which reads out the light from that tile directly. This configuration is optimised for light collection and for automatic assembly with a glue dispenser and a pick-and-place machine.
To better understand the time resolution achievable with the SiPM-on-tile technology, we have performed detailed measurements in beam tests at the DESY II test beam facility [8], probing different scintillator tile sizes. In addition, we introduce measurement methods that allow us to observe the time structure of the scintillation process, the time structure of the light collection in the scintillator tile, and the response of the SiPM and measurement electronics separately.
@@ 1.1 Experimental Equipment
Modules based on a single SiPM-on-tile element were developed for the CALICE T3B experiment [9]. These modules were further developed and used within the CLAWS project for beam background studies and a beam abort system for the SuperKEKB collider [10, 11, 12]. For the measurements discussed in this publication, we used the latest iteration of the CLAWS boards, the phase 3 CLAWS shown in figure 1. Each boards contains a single SiPM under a plastic scintillator tile, as well as the corresponding amplifier electronics.
The SiPMs installed on the CLAWS boards are Hamamatsu S13360-1325PE [13], the same that are used in the AHCAL technological prototype [6]. On each CLAWS module, the signal from the SiPM is amplified by a BGA614 darlington transistor and converted to a differential signal using a LMH6553 differential amplifier.
@@
Up to four CLAWS modules are connected to one receiver box using Cat6A ethernet cables. The cables are used both to supply power to the CLAWS modules and to transmit the differential analog signals from the sensor. The SiPM bias voltage for up to four CLAWS modules is generated in the receiver box and supplied to the front-end via the ethernet cables. LHM6553 differential amplifiers located in the receiver box convert the differential signals to 50 $\Omega$ single ended signals on BNC connectors. The full analog waveforms of each event are recorded with a PicoScope 6804E digital computer-controlled oscilloscope which has a bandwidth of 500 MHz [14].
@@ 1.2 Breaking down the Signal Creation
Figure 2 illustrates the three signal creation steps as well as the different measurement methods. The following section 2 shows our test beam setup and the measured time resolutions. Section 3 discusses the timing properties of BC408 measured with small scintillator cubes, and the laser measurements are presented in section 4. Section 5 contains a discussion of the photon time distributions measured with these different methods.
@@ 2 Test Beam Measurements
Measurements with high-energy charged particles involve all three steps of signal creation (scintillation, light collection and light detection and measurement with the SiPM). We used electrons with an energy of 3 GeV at DESY’s test beam facility [8]. In two test beam weeks of the CALICE AHCAL group, in October 2020 and 2021, we placed our timing setup in front of the AHCAL prototype, such that both experiments used the same electron beam.
@@ 2.1 Experimental Setup
The measurements are controlled by a python script that regularly (approximately every 5 minutes) interrupts for calibration. A calibration run is a separate measurement where each channel is recorded individually, one after another. The trigger level is set to approximately 0.5 p.e which means that it triggers mostly on dark counts where a single pixel of the SiPM fired. Accordingly, the voltage range is set to $\pm 50$ mV. With these calibration measurements the stability of the SiPM gain is monitored for each channel. The integral of the calibration waveforms is used to estimate the number of photoelectrons for the recorded signal waveforms.
@@
$\times 20\times 3~\text{mm}^{3} 20×20$$\times 20\times 3~\text{mm}^{3}$ $\pm 1$ V, 8 bit External October 2021
$\times 30\times 3~\text{mm}^{3} 30×30$$\times 30\times 3~\text{mm}^{3}$ $\pm 1$ V, 8 bit Internal October 2020
@@
Calibration Trigger on 1 p.e. peak $\pm 50$ mV, 8 bit Internal
@@ 2.2 Results: Light Yield
The light yield of a detector module is defined as the most probable number of photoelectrons measured at the SiPM for minimum ionization (1 MIP of energy deposition) in the scintillator tile. The light yield is thus given by the most probable signal area of a minimum ionizing particle signal divided by the area of a 1 p.e. signal.
@@
To extract the most probable energy deposition, we performed a fit to a Landau distribution convolved with a Gaussian distribution, commonly referred to as "Langauss". The Gaussian part accounts for detector noise and measurement uncertainties. The energy deposition in the smallest $20\times 20\times 3~\text{mm}^{3}$ scintillator tiles (figure 4(a)) agrees very well with the expected Langauss behaviour.
The distributions for the larger scintillator tiles (figures 4(b) and 4(c)) on the other hand show an excess of events at energies above the minimum ionization peak [^2] Especially the run with the $40\times 40\times 3~\text{mm}^{3}$ scintillator tiles (figure 4(c)) has a high number of multi-particle events, for once because larger tiles have a higher geometric cross section, and on the other hand because a beam collimator with a larger aperture was used during this measurement.. This can be attributed to events where more than one particle deposits energy in the scintillator tiles. To make sure that no double particle events are included in the analysis, an upper limit for the fit range was set manually before the fit, indicated by the black lines in figure 4. Effectively, these limits correspond to an energy of approximately 1.5 MIP.
@@ 2.3 Results: Time Resolution
Hit times outside of the interval $\pm 5$~ns are rejected since they tend to be due to early hits caused by baseline noise on one of the two channels.
@@
Table 3: Single channel time resolution at minimum ionization $\sigma_{1MIP}$ for different BC408 scintillator tiles measured at the test beam. The total uncertainty includes the uncertainty from the fit as well as expected fluctuations in the time resolution that arise from tile-to-tile differences. The rightmost column is a comparison of the time resolution scaling with the tile area, assuming that the light yield is the only relevant effect.
$\sigma_{1MIP}$ Fit Fit uncertainty Total uncertainty $\sigma_{t}\cdot A^{-0.25}$
Tile size (ns) red. $\chi^{2}$ (ns) (ns) $\text{ns}/\sqrt{\text{cm}}$
$20\times 20\times 3~\text{mm}^{3}$ 0.3847 3.10 0.0004 0.023 $0.272\pm 0.016$
$30\times 30\times 3~\text{mm}^{3}$ 0.5785 1.54 0.0006 0.035 $0.33\pm 0.02$
$40\times 40\times 3~\text{mm}^{3}$ 0.7027 1.76 0.0008 0.042 $0.35\pm 0.02$
@@
As a consequence of equation 2.3, the time resolution of each scintillator tile scales with $\sigma_{t}\propto 1/\sqrt{LY}$. In addition, we know that the light yield scales with $LY\propto 1/\sqrt{A}$. If the light yield of the scintillator tile was the only contribution to the time resolution, then we would expect that $\sigma_{t}\cdot A^{-0.25}$ is constant. However, as shown in table 3, this quantity deviates by more than $3\sigma$ between $20\times 20\times 3~\text{mm}^{3}$ and $30\times 30\times 3~\text{mm}^{3}$. This indicates that there is an additional geometry-dependent contribution to the time resolution, which will be investigated in the following.
@@ 3 Timing Properties of BC408
For measuring the timing properties of BC408, we use a setup that is designed to avoid geometric effects as far as possible and only detects light directly emitted from the scintillator. Scintillator cubes measuring $5\times 5\times 5~\text{mm}^{3}$ are used instead of larger tiles to minimize the path that photons travel. Instead of ESR foil, the cubes are wrapped in black foil to suppress multiple reflections. The scintillator cubes are excited using electrons from a 90Sr source. The signal of one cube is used for the timing analysis, and the second cube is used to provide a coincidence of two cubes as a trigger signal. This ensures full penetration of the first cube, and thus scintillation light emission over the full depth of the scintillator element. The measurement setup is illustrated in figure 7(a), and the measurement settings are summarised in table 4.
@@
Each photon triggers the same pixel response in the SiPM and the shape of this 1 p.e. signal is known from calibration measurements. Therefore, the arrival times of single photons at the SiPM can be estimated from the analog waveform measured with the oscilloscope. To this end we have developed a waveform decomposition algorithm, inspired by the signal reconstruction in the T3B experiment [9]. Starting from the measured waveform, the algorithm iteratively subtracts 1 p.e. signals until only the baseline noise remains. The points in time where the 1 p.e. responses were subtracted are used as photon hit times.
For each event, the photon hit times are corrected by the time of arrival of the particle, which is determined using a 25 % constant fraction on the raw signal. This step is necessary to mitigate the effects of trigger jitter.
A detailed description of the waveform decomposition algorithm, as well as a discussion of its performance, can be found in [17].
The photon emission time distribution of BC408 resulting from the waveform decomposition is shown in figure 7(b). This distribution, which is a material property of BC408, will be discussed together with the light collection process in section 5.
@@
BC408 $5\times 5\times 5~\text{mm}^{3}$ cubes $\pm 500$ mV, 8 bit 0.4 ns Coincidence of both cubes
Laser measurements $\pm 1$ V, 8 bit 0.4 ns Laser driver
(a) Illustration of the measurement setup.
(b) Measured emission time spectrum of BC408
Figure 7: Small scintillator cubes are used to mitigate geometric effects when measuring the photon emission time spectrum of the scintillator material. Left: An illustration of the setup. The scintillator cubes are made of BC408 and have a dimension of $5\times 5\times 5~\text{mm}^{3}$. A black foil shields the cubes from ambient stray light and crosstalk. The oscilloscope is configured as a coincidence trigger on both channels. Right: Photon time spectrum extracted from the measured waveforms using decomposition of the waveform into 1 p.e. peaks.
@@ 4 Laser Measurements
Figure 8 shows an illustration and a photograph of the laser setup. The light intensity that reaches the SiPMs is controlled using a set of neutral density optical filters and is monitored with a power meter. The beam is split and directed to two different SiPMs, which allows us to calculate the time resolution from the hit time difference, as described in section 2.3. The trigger signal is directly provided from the laser driver.
@@
To probe the response of the SiPM and electronics to very short light pulses, the laser beam is aimed directly onto the SiPM. This allows us to estimate the timing uncertainty that arises from the light sensor, the electronics and the data acquisition, i.e. the timing uncertainty of the light detection.
@@
5 Discussion of Photon Time Distributions
5.1 Light Detection and Hardware Response
Figure 9(a) shows the time resolution of the light detection which was obtained from laser measurements, compared to the time resolution of a $30\times 30\times 3~\text{mm}^{3}$ scintillator tile at the test beam. If we assume that the timing uncertainty from the electronics and SiPM is uncorrelated to the effects of the scintillator tile (scintillation and light collection), we can write the overall time resolution of the SiPM-on-tile system as
@@ 5.1 Light Detection and Hardware Response
The green data points in figure 9(a) are the time resolution of the test beam measurement without the uncertainties induced by the time resolution of the SiPM:
@@
This can also be seen the time distribution of photons arriving at the SiPM, shown in figure 9(b). Compared to the time distribution of scintillation photons, the laser measurement yields a very short time distribution. Further discussion on this plot can be found in [17].
(a) Energy dependent time resolution
(b) Photoelectron time distribution
Figure 9: Comparison of the test beam measurements (blue) and a pulsed laser signal directly on the SiPM (orange). The left plot compares these two measurements in terms of energy-dependent time resolution. The green curve is the computed time resolution of the test beam measurement without the timing uncertainties of the SiPM. The common energy scale is photoelectrons, and for reference the MIP scale of the test beam measurement is shown at the top axis. The right plot compares the photoelectron time distribution of test beam and laser-on-SiPM measurements. The photon hit times were extracted by decomposition of the waveform into 1 p.e. peaks (see section 3).
@@
All in all, the laser measurements show that the SiPM, CLAWS, amplifier electronics and oscilloscope together are still significantly faster than other signal components. Therefore, the contribution of the hardware to the time structure of the measured signals can be neglected.
5.2 Scintillation and Light Collection